### v1

# Linearization in magnetoelasticity

Abstract.
Starting from a model of nonlinear magnetoelasticity where magnetization is defined in the Eulerian configuration while elasticity is in the Lagrangean one, we rigorously derive a linearized model that coincides with the standard one that already appeared in the literature, see, e.g., [12] and where the zero-stress strain is quadratic in the magnetization. The relation of the nonlinear and linear model is stated in terms of the $\Gamma$-convergence and convergence of minimizers.
Key words and phrases:
Linearization, magnetoelasticity
1. Introduction
Magnetoelasticity, also known as magnetostriction, is a characteristic observed in certain solids wherein there is a significant interdependence between mechanical and magnetic properties. This coupling manifests as reversible mechanical deformations that can be induced by applying an external magnetic field. This phenomenon holds considerable practical value, especially in the design of sensors, actuators, and various innovative functional-material devices, see e.g.  [17].
The origin of magnetoelasticity can be traced to the intricate interplay between the crystallographic structure of the material, where different crystals exhibit distinct easy axes of magnetization, and magnetic domains. In the absence of external magnetic fields, these magnetic domains align to minimize long-range dipolar effects, leading to a generally minimal or even negligible magnetization of the medium. Upon the application of an external magnetic field, the magnetic domains undergo a reorientation towards it. This reorientation leads to the emergence of a macroscopic deformation since magnetizations are linked to specific stress-free reference strains. As the magnetic field’s intensity increases, an increasing number of magnetic domains align themselves so that their principal axes of anisotropy align with the magnetic field in each region, eventually reaching saturation. For a deeper exploration of the foundational aspects of magnetoelasticity, refer to [7, 13].
Magnetoelasticity is a dynamic area of mathematical research reflecting modeling in linear and nonlinear regimes [12, 24] and analysis in the static and quasistatic setting [2, 5, 4] but also coupled with thermal and other phenomena, e.g.  [21]. Moreover, dimension-reduction problems covering lower-dimensional structures in various scaling regimes were recently analyzed, too; see, for instance, [4] or [11].
The aim of our contribution is to establish a relationship between nonlinear and linear models of magnetoelasticity in the static case. To our best knowledge, it is the first situation where this program is undertaken in the combined Eulerian and Lagrangian description.
Let $\Omega$ be a (connected) bounded domain in $\mathbb{R}^{d}$ $(d\geq 2)$ with Lipschitz boundary $\partial\Omega$, representing a reference configuration occupied by an elastic body. We fix $\Gamma$ open subset of $\partial\Omega$ such that $\mathcal{H}^{d-1}(\Gamma)>0$, where $\mathcal{H}^{d-1}$ denotes the $(d-1)$-dimensional Hausdorff measure. We consider the magnetoelastic energy of the following form [7, 18, 5, 23]:
(1.1) $\mathcal{G}(y,m):=\int_{\Omega}W(\nabla y,m\circ y){\,\rm d}x+\frac{1}{2}\int_{y(\Omega)}|\nabla m|^{2}{\,\rm d}y+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m}|^{2}{\,\rm d}z$
for a deformation $y\in W^{1,p}(\Omega;\mathbb{R}^{d})$, $y=y_{0}$ on $\Gamma$, and a magnetization $m\in W^{1,2}(y(\Omega);\mathbb{R}^{d})$. In (1.1), the first term represents the elastic energy, the second term is the exchange energy, and the last term stands for magnetostatic energy; $\mu_{0}$ is the permittivity of void and $v_{m}$ is the magnetostatic potential generated by $m$. In particular, $v_{m}$ is a solution to the Maxwell equation $\nabla\cdot(-\mu_{0}\nabla v_{m}+\chi_{y(\Omega)}m)=0$ in $\mathbb{R}^{d}$ with $\chi_{y(\Omega)}$ the characteristic function of $y(\Omega)$. Furthermore, we assume that the specimen temperature is fixed and therefore the Heisenberg normalized constraint
(1.2) $|m\circ y|\,\det\nabla y=1\text{ a.e.~{}in }\Omega$
holds for all deformations.
As in [14], we suppose the deformation gradient $F$ can be multiplicatively decomposed into magnetic and elastic parts $F=F_{m}F_{el}$, where the magnetic part $F_{m}$ is volume-preserving, i.e., $\det F_{m}=1$. In our setting, we assume $F_{m}:=\exp(-E(\mathcal{M}))$, i.e., $F^{-1}_{m}=\exp(E(\mathcal{M}))$, where $E(\mathcal{M})$ is defined as
$E(\mathcal{M}):=-\mathcal{M}\otimes\mathcal{M}+\frac{1}{d}I\,.$
Here, $\mathcal{M}\in\mathbb{R}^{d}$ is a placeholder for the Lagrangean magnetization whose length $|\mathcal{M}|$ satisfies $\mathcal{M}\cdot\mathcal{M}=|\mathcal{M}|^{2}=1$, and $I$ is the identity matrix in $\mathbb{R}^{d\times d}$. Note that $\operatorname{tr}E(\mathcal{M})=0$, so that, indeed, $\det F_{m}^{-1}=1$. In view of the Heisenberg constraint (1.2), $M(x)=m(y(x))\det\nabla y(x)$ satisfies $|M(x)|=1$ for $x\in\Omega$. Consequently, if $\omega\subset\Omega$ we formally have the change-of-variables formula implying
$\int_{\omega}M(x)\,{\,\rm d}x=\int_{y(\omega)}m(z)\,{\,\rm d}z,$
and $M$ and $m$ transform as volume densities in the reference (Lagrangean) and deformed (Eulerian) configurations, respectively.
Then we define
$e(F,m)=E(M)=-M\otimes M+\frac{1}{d}I=-(\det F)^{2}m\otimes m+\frac{1}{d}I\ .$
In the case $F=I$, we simplify the notation by setting $e(I,m)=e(m):=-m\otimes m+\frac{1}{d}I$. Further, we assume that the elastic energy density is of form
$W(F,m)=\Phi\left(\exp(e(F,m))F\right)=\Phi(F_{el})\ ,$
where $c>0$ is a material constant. We recall that the matrix exponential $\exp(A)=\sum_{k=0}^{\infty}A^{k}/k!$ is defined for every $A\in\mathbb{R}^{d\times d}$, setting also $A^{0}=I$. Here, $\Phi$ is a nonnegative function that is minimized at the identity.
In the small-strain regime, we set $y_{\varepsilon}=id+\varepsilon u$ and assume that the influence of the magnetization on the elastic deformation of the specimen vanishes along $\varepsilon\to 0$ in the sense that $F_{m,\varepsilon}=\exp(\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varepsilon E(M))$ and $\lim_{\varepsilon\to 0}F_{m,\varepsilon}=I$. Then, we consider the rescaled energy
(1.3) $\mathcal{G}_{\varepsilon}(u,m)$ $:=\frac{1}{\varepsilon^{2}}\int_{\Omega}W(Id+\varepsilon\nabla u(x),m(x+\varepsilon u(x))){\,\rm d}x+\frac{1}{2}\int_{y_{\varepsilon}(\Omega)}|\nabla m(y)|^{2}{\,\rm d}y+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m}(z)|^{2}{\,\rm d}z$
$=\frac{1}{\varepsilon^{2}}\Phi(\exp(\varepsilon E(M))(I+\varepsilon\nabla u(x)))\,{\,\rm d}x\frac{1}{2}\int_{y_{\varepsilon}(\Omega)}|\nabla m(y)|^{2}{\,\rm d}y+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m}(z)|^{2}{\,\rm d}z.$
We formally get the expression for the elastic strain
$F_{el,\varepsilon}=F^{-1}_{m,\varepsilon}\nabla y_{\varepsilon}=\operatorname{exp}(\varepsilon E(M))\cdot(I+\varepsilon\nabla u)\approx I+\varepsilon(\nabla u+e(m)).$
Moreover, a formal Taylor expansion shows that the limit energy is, as in [12], of the form:
(1.4) $\mathcal{G}_{0}(u,m):=\frac{1}{{2}}\int_{\Omega}\mathbb{C}(\bm{\epsilon}(u)+e(m)):(\bm{\epsilon}(u)+e(m)){\,\rm d}x+\frac{1}{2}\int_{\Omega}|\nabla m|^{2}{\,\rm d}x+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m}|^{2}{\,\rm d}z,$
where $e(u):=\frac{1}{2}(\nabla u^{T}+\nabla u)$ is the symmetric part of the displacement gradient and $\mathbb{C}:=D^{2}\Phi(I)$ is the elasticity tensor. Note that (1.4) suggests a decomposition of the total strain $\bm{\epsilon}(u)$ into the magnetic part $-e(m)=m\otimes m-I/d$ and the elastic part $\bm{\epsilon}(u)+e(m)$. This decomposition was already used, e.g., in [12, 19, 20]. The aim of this contribution is to derive (1.4) rigorously by means of $\Gamma$-convergence.
We use the following assumptions for $p>d$:
  1. (a)
    $\Phi$ is frame-indifferent;
  2. (b)
    $\Phi$ belongs to $C^{2}$ in some neighbourhood of $SO(d)$;
  3. (c)
    $\Phi(F)=0$ if $F\in SO(d)$;
  4. (d)
    $\Phi(F)\geq g_{p}(\operatorname{dist}(F,SO(d)))$ for some $p$, where
    (1.5) $g_{p}(t):=\begin{cases}\frac{t^{2}}{2},&\text{if }0\leq t\leq 1,\\ \frac{t^{p}}{p}+\frac{1}{2}-\frac{1}{p},&\text{if }t>1.\end{cases}$
  5. (e)
    there exist $a>1$ and $C>0$ such that
    $\Phi(F)\geq C\bigg{(}\frac{1}{(\det F)^{a}}-1\bigg{)}\qquad\text{for every $F\in\mathbb{M}^{d}$.}$
Frame-indifference (a) implies that $W(F,m)=W(F,-m)=W(QF,Qm)$ for all $F\in\mathbb{R}^{d\times d}$ all $Q\in\textrm{SO}(d)$, and all $m\in\mathbb{R}^{d}$. Moreover, for every $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in\mathbb{R}^{d\times d}$
(1.6) $D^{2}\Phi(I)[\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F,F\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}]=D^{2}\Phi(I)[\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{sym},F_{sym}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}].$
Together with (d), it implies that $D^{2}\Phi(I)[\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F,F\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}]=0$ if $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F=-F^{\top}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$ and
(1.7) $D^{2}\Phi(I)[\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{sym},F_{sym}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}]\geq|\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F_{sym}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|^{2}.$
Above $F_{sym}$ stands for the symmetric part of $F$. While (b) together with $\Phi(I)=D\Phi(I)=0$ leads to
(1.8) $\Phi(I+F)\geq\frac{1}{2}D^{2}\Phi(I)[F,F]-\eta(|F|)|F|^{2},$
where $\eta$ is an increasing nonnegative function such that $\eta(t)\leq C_{R}t$ for $|t|\leq R$ and $R>0$.
In what follows, the work of a mechanical load is modeled by a continuous linear functional $\mathscr{L}\colon L^{2}(\Omega;\mathbb{R}^{d})\to\mathbb{R}$ and the magnetic loading is described by a functional $\mathscr{M}\colon L^{2}(\Omega;\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})\to\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}$
$\mathscr{L}(y):=\int_{\Omega}f\cdot y{\,\rm d}x,\qquad\text{and}\qquad\mathscr{M}(y,m):=\int_{y(\Omega)}h\cdot m{\,\rm d}z$
where $f\in L^{2}(\Omega;\mathbb{R}^{d})$ is a body force and $h\in L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$ represents an external magnetic field, while $\mathscr{M}$ is called the Zeeman energy.
If $y\in W^{1,p}(\Omega;\mathbb{R}^{d})$ represents the deformation of the elastic body, the stable equilibria of the elastic body are obtained by minimizing the functional
$\int_{\Omega}W(\nabla y,m\circ y){\,\rm d}x-\mathscr{L}(y)-\mathscr{M}(y,m),$
under the prescribed boundary conditions. We are interested in the case where the load has the form $\varepsilon\mathscr{L}$, and we want to study the behavior of the solution as $\varepsilon$ tends to zero. We write
$y=x+\varepsilon u$
and we assume the Dirichlet boundary condition of the form
$v=x+\varepsilon w\quad\mathcal{H}^{n-1}\text{-a.e. on }\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma,\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$
with a given function $w\in W^{2,\infty}(\Omega;\mathbb{R}^{d})$. Denoting a set of functions satisfying boundary conditions as
$W^{1,2}_{w}(\Omega;\mathbb{R}^{d}):=\left\{u\in W^{1,2}(\Omega;\mathbb{R}^{d}):u=w\quad\mathcal{H}^{n-1}\text{-a.e. on }\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\right\}\!,$
the corresponding minimum problem for $u$ becomes
$\min\biggl{\{}\int_{\Omega}W(I+\varepsilon\nabla u(x),m(x+\varepsilon u(x))){\,\rm d}x-\varepsilon\mathscr{L}(\varepsilon u)$ $-\mathscr{M}(id+\varepsilon u,m):$
$(u,m)\in W^{1,p}_{w}(\Omega;\mathbb{R}^{d})\times W^{1,2}(y_{\varepsilon}(\Omega);\mathbb{R}^{d})\biggr{\}}\!.$
The class of admissible deformations is given by
(1.9) $\mathcal{Y}:=\left\{y\in W^{1,p}(\Omega;\mathbb{R}^{d}):\det\nabla y>0\text{ a.e. in }\Omega,\>y\text{ is a.e.~{}injective in }\Omega\right\}\!.$
For $w\in W^{2,\infty}(\Omega;\mathbb{R}^{d})$ and $\varepsilon>0$, define the class of admissible states as
$\mathcal{A}^{w}_{\varepsilon}:=\bigl{\{}(u,m)$ $\in W^{1,2}_{w}(\Omega;\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d};\mathbb{R}^{d}):y_{\varepsilon}:=id+\varepsilon u\in\mathcal{Y},$
$\chi_{y_{\varepsilon}(\Omega)}m\in W^{1,2}(y_{\varepsilon}(\Omega),\mathbb{R}^{d}),\>|m\circ y_{\varepsilon}|\,\det\nabla y_{\varepsilon}=1\text{ a.e.~{}in }\Omega\bigr{\}},$
and
$\mathcal{A}^{w}_{0}:=W^{1,2}_{w}(\Omega;\mathbb{R}^{d})\times W^{1,2}(\Omega;\mathbb{S}^{d-1}).$
The main result of this work is the following $\Gamma$-convergence statement.
Theorem 1.1.
Assume that $W\colon\mathbb{M}^{d}\times\mathbb{R}^{d}\rightarrow[0,\infty]$ satisfies conditions (a)(e) for some $p>d$, and let $w\in W^{2,\infty}(\Omega;\mathbb{R}^{d})$. For every $\varepsilon_{j}\to 0$ we have that
$\mathcal{G}_{\varepsilon_{j}}\stackrel{{\scriptstyle\Gamma}}{{\longrightarrow}}\mathcal{G}_{0},\quad\text{ as }j\to\infty$
in the weak topology of $W^{1,2}(\Omega;\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$.
Theorem 1.1 is completed by the following theorem, stating the convergence of minima and minimizers. Here, we consider loading terms as well. In particular, for every $\varepsilon>0$ and every $(u,m)\in\mathcal{A}^{w}_{\varepsilon}$ we set
$\mathcal{F}_{\varepsilon}(u,m):=\mathcal{G}_{\varepsilon}(u,m)-\mathscr{L}(u)-\mathscr{M}(id+\varepsilon u,m).$
Theorem 1.2.
Under the hypotheses of Theorem 1.1, for every $\varepsilon>0$ define $s_{\varepsilon}:=\inf\{\mathcal{F}_{\varepsilon}(u,m):(u,m)\in\mathcal{A}^{w}_{\varepsilon}\}$ and let $\left\{(u_{\varepsilon},m_{\varepsilon})\right\}_{\varepsilon>0}$ be a sequence such that
$\mathcal{F}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})=s_{\varepsilon}+o(1).$
Then, $u_{\varepsilon}\rightharpoonup u_{0}$ weakly in $W^{1,2}(\Omega;\mathbb{R}^{d})$, $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m_{0}$ in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}\rightharpoonup\chi_{\Omega}\nabla m_{0}$ weakly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, where $(u_{0},m_{0})\in\mathcal{A}_{0}^{w}$ is the solution to the minimum problem
$s_{0}:=\min\bigl{\{}\mathcal{F}_{0}(u,m):(u,m)\in\mathcal{A}^{w}_{0}\bigr{\}},$
where $\mathcal{F}_{0}(u,m):=\mathcal{G}_{0}(u,m)-\mathscr{L}(u)-\mathscr{M}(id,m)$ and $\mathcal{G}_{0}$ is defined by (1.4). Moreover, $s_{\varepsilon}\rightarrow s_{0}$.
Theorem 1.3.
Under the hypotheses of Theorem 1.1, $(u_{\varepsilon},m_{\varepsilon})\to(u_{0},m_{0})$ strongly in $W^{1,2}(\Omega;\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$.
The proof of Theorems 1.11.3 is postponed to Sections 4. Compactness issues are discussed in Section 3.
2. Auxiliary results
Define a homogeneous Sobolev space $L^{1,2}(\mathbb{R}^{d})$ as the space of $\varphi\in L^{2}_{loc}(\mathbb{R}^{d})$ such that $\nabla\varphi\in L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. Then magnetic potential $v_{m}\in L^{1,2}(\mathbb{R}^{d})$ generated by magnetization $m$ exists, unique up to a constant, and continuously depends on $m$.
Proposition 2.1.
For every $f\in L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$ there exists $v_{f}\in L^{1,2}(\mathbb{R}^{d})$ which is unique up to additive constants, such that
$\int_{\mathbb{R}^{d}}(-\nabla v_{f}+f)\cdot\nabla\varphi{\,\rm d}z=0\qquad\text{for all }\varphi\in L^{1,2}(\mathbb{R}^{d})$
and
$\|\nabla v_{f}\|_{L^{2}}\leq\|f\|_{L^{2}}.$
Moreover, if $f_{j}\to f$ strongly in $L^{2}$, then
$\|\nabla v_{f_{j}}\|_{L^{2}}\to\|\nabla v_{f}\|_{L^{2}}.$
Proof.
This proposition is classical, see for example [2, Proposition 8.8 & Theorem 8.9] or [23, Proof of Theorem 4.1]. ∎
Let us now recall sufficient conditions for a deformation to be injective if the gradient of the displacement is small enough in the operator norm, i.e., $|B|_{O}=\sup_{v\in\mathbb{R}^{d}}\frac{|Bv|}{|v|}$.
Theorem 2.2 ([8, Theorem 5.5-1]).
Let $\Omega\subset\mathbb{R}^{d}$ be an open set.
  1. (1)
    Let $y=id+u\colon\Omega\to\mathbb{R}^{d}$ be differentiable in $x\in\Omega$. Then
    $|\nabla u|_{O}<1\Rightarrow\det\nabla y>0.$
  2. (2)
    Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary. Then there exists a constant $c(\Omega)$ such that any $y=id+u\in C^{1}(\bar{\Omega};\mathbb{R}^{d})$ with $\sup_{x\in\bar{\Omega}}|\nabla u(x)|_{O}<c(\Omega)$ is injective.
To establish compactness, we need the following geometric rigidity, proven in [15, Theorem 3.1] for $p=2$ and in [9, Section 2.4] for $p\in(1,+\infty)$.
Theorem 2.3.
There exists a constant $C=C(\Omega,p)>0$ such that for every $v\in W^{1,p}(\Omega;\mathbb{R}^{d})$ there exists a constant rotation $R\in SO(d)$ satisfying
(2.1) $\int_{\Omega}|\nabla v-R|^{p}{\,\rm d}x\leq C\int_{\Omega}g_{p}({\operatorname{dist}}(\nabla v,SO(d))){\,\rm d}x,$
where $g_{p}$ is defined by (1.5). In particular, we may choose $R$ such that
(2.2) $\left|R-\frac{1}{\mathcal{L}^{d}(\Omega)}\int_{\Omega}\nabla v\,{\,\rm d}x\right|={\operatorname{dist}}\left(\frac{1}{\mathcal{L}^{d}(\Omega)}\int_{\Omega}\nabla v\,{\,\rm d}x,SO(d)\right)\!.$
Proof.
It is enough to resort to the classical rigidity estimates for $p\in(1,+\infty)$ and notice that
$\frac{1}{2p}(t^{p}+t^{2})\leq\frac{1}{p}\max\{t^{p},t^{2}\}\leq g_{p}(t)\leq\frac{1}{2}\max\{t^{p},t^{2}\}\leq\frac{1}{2}(t^{p}+t^{2}).$
The particular choice of $R$ as in (2.2) can be ensured arguing, e.g., as in [1, Theorem 3.1]. ∎
3. Compactness
This section is devoted to the proof of the following compactness result for sequences with bounded energy $\mathcal{G}_{\varepsilon}$.
Proposition 3.1 (Compactness).
Under the assumptions of Theorem 1.2, let $(u_{\varepsilon},m_{\varepsilon})\in\mathcal{A}^{w}_{\varepsilon}$ be such that, for $y_{\varepsilon}:=id+\varepsilon u_{\varepsilon}$, it holds
(3.1) $\sup_{\varepsilon>0}\,\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})<+\infty.$
Then, there exist $(u,m)\in W^{1,2}_{w}(\Omega;\mathbb{R}^{d})\times W^{1,2}(\Omega;\mathbb{S}^{d-1})$ such that, up to a subsequence,
[…]
Before proving Proposition 3.1, we show the following lemmas, which provide boundedness of the displacement variable in terms of the energy $\mathcal{G}_{\varepsilon}$.
Lemma 3.2.
Under the assumptions of Theorem 1.1, there exists a positive constant $C=C(\Omega,d,p)>0$ such that for every $\varepsilon\in(0,1)$ and every $(u,m)\in\mathcal{A}^{w}_{\varepsilon}$ it holds
(3.2) $\varepsilon^{2}\mathcal{G}_{\varepsilon}(u,m)\geq C\int_{\Omega}g_{p}(\operatorname{dist}(\nabla y_{\varepsilon},SO(d))\,{\,\rm d}x-C\varepsilon^{2},$
for $y_{\varepsilon}=id+\varepsilon u$.
Proof.
Let $(u,m)\in\mathcal{A}^{w}_{\varepsilon}$ and $\varepsilon\in(0,1)$. By assumption (d) we can estimate
(3.3) $\varepsilon^{2}\mathcal{G}_{\varepsilon}(u,m)\geq C\int_{\Omega}g_{p}(\operatorname{dist}\big{(}\exp(\varepsilon e(\nabla y_{\varepsilon},m\circ y_{\varepsilon})\nabla y_{\varepsilon};SO(d)\big{)}\,{\,\rm d}x.$
By definition of $e(\nabla y_{\varepsilon},m\circ y_{\varepsilon})$, by the properties of the exponential map, and by the fact that $\det{\nabla y_{\varepsilon}}|m\circ y_{\varepsilon}|=1$ in $\Omega$, we have that for $x\in\Omega$ and $R\in SO(d)$
$\operatorname{dist}$ $(\nabla y_{\varepsilon};\exp(-\varepsilon e(\nabla y_{\varepsilon},m\circ y_{\varepsilon}))SO(d))\leq|\nabla y_{\varepsilon}-\exp(-\varepsilon e(\nabla y_{\varepsilon},m\circ y_{\varepsilon}))R|$
$\leq|\exp(-\varepsilon e(\nabla y_{\varepsilon},m\circ y_{\varepsilon}))|\,|\exp(\varepsilon e(\nabla y_{\varepsilon},m\circ y_{\varepsilon}))\nabla y_{\varepsilon}-R|\leq C|\exp(\varepsilon e(\nabla y_{\varepsilon},m\circ y_{\varepsilon}))\nabla y_{\varepsilon}-R|.$
This implies that for a.e. $x\in\Omega$
$\operatorname{dist}(\nabla y_{\varepsilon};\exp(-\varepsilon e(\nabla y_{\varepsilon},m\circ y_{\varepsilon}))SO(d))\leq C\,\operatorname{dist}(\exp(\varepsilon e(\nabla y_{\varepsilon},m\circ y_{\varepsilon})\nabla y_{\varepsilon};SO(d)))\,.$
Again by the properties of the exponential, we further estimate
(3.4) $\operatorname{dist}(\nabla y_{\varepsilon};SO(d))\leq C\,\operatorname{dist}(\nabla y_{\varepsilon};\exp(-\varepsilon e(\nabla y_{\varepsilon},m\circ y_{\varepsilon})SO(d))+C\varepsilon\qquad\text{a.e.~{}in~{}$\Omega$.}$
Combining (3.3)–(3.4) we get (3.2). ∎
Lemma 3.3.
Under the assumptions of Theorem 1.1, there exists a positive constant $C=C(\Omega,d,p)>0$ such that for every $\varepsilon>0$, every $(u,m)\in\mathcal{A}^{w}_{\varepsilon}$, and every $\varepsilon>0$ the following estimates hold:
(3.5) $\int_{\Omega}|\nabla u|^{2}{\,\rm d}x\leq C\Bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\Gamma}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
(3.6) $\int_{\Omega}|\varepsilon\nabla u|^{p}{\,\rm d}x\leq C\varepsilon^{2}\bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\Gamma}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)}.$
Proof.
Along the proof, we denote by $C$ a generic positive constant depending only on $\Omega$, $d$, and $p$, but not on $\varepsilon$ and on $(u,m)\in\mathcal{A}_{\varepsilon}^{w}$.
Let us prove (3.5). For $(u,m)\in\mathcal{A}^{w}_{\varepsilon}$ we set $y_{\varepsilon}:=id+\varepsilon u$. Applying the Geometric Rigidity Theorem 2.3 and Lemma 3.2, we deduce that there exists $R_{\varepsilon}\in SO(d)$ such that
$\int_{\Omega}|\nabla y_{\varepsilon}-R_{\varepsilon}|^{2}{\,\rm d}x\leq C\int_{\Omega}g_{p}(\operatorname{dist}(\nabla y_{\varepsilon};SO(d))){\,\rm d}x\leq C\varepsilon^{2}\big{(}\mathcal{G}_{\varepsilon}(u,m)+1).$
By triangle inequality, we continue with
(3.7) $\int_{\Omega}|\nabla y_{\varepsilon}-I|^{2}{\,\rm d}x\leq C\big{(}\varepsilon^{2}\big{(}\mathcal{G}_{\varepsilon}(u,m)+1)+|R_{\varepsilon}-I|^{2}\big{)}.$
Arguing as in [10, Proposition 3.4], we get that
(3.8) $|R_{\varepsilon}-I|^{2}\leq C\varepsilon^{2}\bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\Gamma}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
which, together with (3.7), implies (3.5).
As for (3.6), we notice that, by Theorem 2.3 and Lemma 3.2, for the same rotation $R_{\varepsilon}$ it holds
$\int_{\Omega}|\nabla y_{\varepsilon}-R_{\varepsilon}|^{p}{\,\rm d}x\leq C\int_{\Omega}g_{p}(\operatorname{dist}(\nabla y_{\varepsilon};SO(d))){\,\rm d}x+C\varepsilon^{2}\leq C\varepsilon^{2}\big{(}\mathcal{G}_{\varepsilon}(u,m)+1\big{)}.$
Hence, recalling (3.8) we deduce that
$\int_{\Omega}|\varepsilon\nabla u|^{p}{\,\rm d}x$ $\leq 2^{p-1}\bigg{(}\int_{\Omega}|\nabla y_{\varepsilon}-R_{\varepsilon}|^{p}{\,\rm d}x+|R_{\varepsilon}-I|^{p}\bigg{)}\leq C\bigg{(}\varepsilon^{2}\mathcal{G}_{\varepsilon}(u,m)+\varepsilon^{2}+|R_{\varepsilon}-I|^{p}\bigg{)}$
$\leq C\bigg{(}\varepsilon^{2}\mathcal{G}_{\varepsilon}(u,m)+\varepsilon^{2}+|R_{\varepsilon}-I|^{2}\bigg{)}\leq C\varepsilon^{2}\bigg{(}\mathcal{G}_{\varepsilon}(u,m)+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\Gamma}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
where we have used the fact that $p>d\geq 2$ and that $R_{\varepsilon}\in SO(d)$, so that $|R_{\varepsilon}-I|\leq 2\sqrt{d}$. This concludes the proof of the proposition. ∎
We are now ready to prove the compactness result of Proposition 3.1.
Proof of Proposition 3.1.
For simplicity of notation, let us set $\Lambda_{\varepsilon}:=\exp(\varepsilon e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}))\nabla y_{\varepsilon}$. Then, by the assumptions on $\Phi$ we have that
$\int_{\Omega}\frac{1}{(\det\nabla y_{\varepsilon})^{a}}{\,\rm d}x$ $=\int_{\Omega}\frac{1}{(\det\Lambda_{\varepsilon})^{a}}{\,\rm d}x$
$\leq C\int_{\Omega}W(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}){\,\rm d}x+C\mathcal{L}^{d}(\Omega)\leq C\varepsilon^{2}\mathcal{G}_{\varepsilon}(y_{\varepsilon},m_{\varepsilon})+C\mathcal{L}^{d}(\Omega).$
Thus, $1/\det\nabla y_{\varepsilon}$ is equi-integrable in $\Omega$.
Furthermore, we recall that (3.6) implies that
(3.9) $\|\nabla y_{\varepsilon}-I\|_{L^{p}}\leq C\varepsilon^{\frac{2}{p}}$
for some positive constant $C$ independent of $\varepsilon$.
For $\delta>0$ let us set
$\Omega_{\delta}:=\{x\in\Omega:\,\operatorname{dist}(x,\partial\Omega)>\delta\}\qquad\text{ and }\qquad\mu_{\varepsilon}^{\delta}:=\chi_{\Omega_{\delta}}m_{\varepsilon}.$
Without loss of generality, we may assume that $\Omega_{\delta}$ is a Lipschitz set. Since $y_{\varepsilon}\to id$ in $W^{1,p}(\Omega;\mathbb{R}^{d})$ and $p>d$, we have that $y_{\varepsilon}\to id$ in $C^{0}(\overline{\Omega};\mathbb{R}^{d})$. Hence, for $\varepsilon>0$ small enough, we may assume that $\Omega_{\delta}\subseteq y_{\varepsilon}(\Omega)$. By change of variables, using the equality $|m_{\varepsilon}\circ y_{\varepsilon}|\det\nabla y_{\varepsilon}=1$ a.e. in $\Omega$, we get that
(3.10) $\int_{\mathbb{R}^{d}}|\mu_{\varepsilon}^{\delta}|^{2}{\,\rm d}x$ $=\int_{\Omega_{\delta}}|m_{\varepsilon}|^{2}{\,\rm d}x$
$\leq\int_{y_{\varepsilon}(\Omega)}|m_{\varepsilon}|^{2}{\,\rm d}x=\int_{\Omega}|m_{\varepsilon}\circ y_{\varepsilon}|^{2}\,\det\nabla y_{\varepsilon}{\,\rm d}x=\int_{\Omega}\frac{1}{\det\nabla y_{\varepsilon}}{\,\rm d}x.$
Moreover, by definition of $\mu_{\varepsilon}^{\delta}$ it holds
(3.11) $\int_{\Omega_{\delta}}|\nabla\mu_{\varepsilon}^{\delta}|^{2}{\,\rm d}x\leq\int_{y_{\varepsilon}(\Omega)}|\nabla m_{\varepsilon}|^{2}{\,\rm d}x\leq\mathcal{G}_{\varepsilon}(y_{\varepsilon},m_{\varepsilon}).$
Therefore, $\mu_{\varepsilon}^{\delta}$ is bounded in $W^{1,2}(\Omega_{\delta};\mathbb{R}^{d})$ uniformly for all $\delta>0$. By a diagonal argument, we find $m\in W^{1,2}_{loc}(\Omega;\mathbb{R}^{d})$ such that $\mu_{\varepsilon}^{\delta_{\varepsilon}}\rightharpoonup m$ weakly in $W^{1,2}_{loc}(\Omega;\mathbb{R}^{d})$. While the Sobolev Embedding Theorems provide that
(3.12) $m_{\varepsilon}(x)=\mu_{\varepsilon}^{\delta}(x)\to m(x)\qquad\text{for a.e.~{}}x\in\Omega_{\delta}.$
Passing to the liminf by $\varepsilon$ and then to the limit by $\delta$ in (3.10) and (3.11), we further deduce that $m\in W^{1,2}(\Omega;\mathbb{R}^{d})$. Note further that the second line in (3.10) ensures that $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}$ is bounded in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and thus converges weakly to some $\mu\in L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. Using pointwise convergence (3.12), we conclude $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}(x)\to\chi_{\Omega}m(x)$ a.e. in $\mathbb{R}^{d}$. Therefore $\mu=\chi_{\Omega}m$ a.e., and, up to a subsequence,
(3.13) $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\rightharpoonup\chi_{\Omega}m\quad\text{ weakly in }L^{2}(\mathbb{R}^{d};\mathbb{R}^{d}).$
Further, applying the Lebesgue–Vitali Convergence Theorem (see [3, Th. 4.5.4] or [22, Ch. 6, Ex. 10]) to $|\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}|^{2}$, we obtain convergence of norms
$\|\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\|_{L^{2}}\to\|\chi_{\Omega}m\|_{L^{2}}.$
This, together with weak convergence (3.13) guarantees strong convergence $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m$ in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. Since the weak convergence in $L^{2}(\Omega;\mathbb{R}^{d})$ of $m_{\varepsilon}\chi_{y_{\varepsilon}(\Omega)}$ to $m$ has been already shown, we deduce the strong $L^{2}$-convergence.
By (3.1) we know that $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}$ is bounded in $L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$. For every test $\varphi\in L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$ we consider
$\int_{\mathbb{R}^{d}}$ $(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m):\varphi{\,\rm d}x$
$=\int_{\Omega_{\delta}}(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m):\varphi{\,\rm d}x+\int_{\mathbb{R}^{d}\setminus\Omega_{\delta}}(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m):\varphi{\,\rm d}x.$
Then, the first integral tends to 00 since $\Omega_{\delta}\subseteq y_{\varepsilon}(\Omega)\cap\Omega$, $m_{\varepsilon}=\mu_{\varepsilon}^{\delta}$ in $\Omega_{\delta}$, and $\mu_{\varepsilon}^{\delta}\rightharpoonup m$ weakly in $W^{1,2}(\Omega_{\delta};\mathbb{R}^{d})$. As for the second integral, we have that
$\bigg{|}\int_{\mathbb{R}^{d}\setminus\Omega_{\delta}}(\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}-\chi_{\Omega}\nabla m):\varphi{\,\rm d}x\bigg{|}\leq C\|\varphi\|_{L^{2}(C_{\delta})},$
where $C_{\delta}=\{x\in\mathbb{R}^{d}:\,\operatorname{dist}(x,\partial\Omega)<\delta\}$. Since $\mathcal{L}^{d}(C_{\delta})\to 0$ as $\delta\to 0$, we infer that $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}$ converges to $\chi_{\Omega}\nabla m$ weakly in $L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$.
We now show that $m_{\varepsilon}\circ y_{\varepsilon}$ converges to $m$ in $L^{1}(\Omega;\mathbb{R}^{d})$. For $\delta>0$ and $\varepsilon\ll 1$ we write
(3.14) $\int_{\Omega_{\delta}}|m_{\varepsilon}\circ y_{\varepsilon}-m|{\,\rm d}x\leq\int_{\Omega_{\delta}}|m_{\varepsilon}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x+\int_{\Omega_{\delta}}|\mu_{\varepsilon}^{\delta}-m|{\,\rm d}x.$
For the second integral on the right-hand side of (3.14) we simply have that $\mu_{\varepsilon}^{\delta}\to m$ in $L^{2}(\Omega_{\delta};\mathbb{R}^{d})$. As for the first term on the right-hand side of (3.14), we notice that $m_{\varepsilon}\circ y_{\varepsilon}$ and $\mu_{\varepsilon}^{\delta}$ are equi-integrable in $\Omega_{\delta}$. Thus, for $\eta>0$ there exists $\rho>0$ such that for every $A\subseteq\Omega_{\delta}$ measurable
$\mathcal{L}^{d}(A)<\rho\ \Rightarrow\ \limsup_{\varepsilon\to 0}\int_{A}|m_{\varepsilon}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x<\eta.$
Let us consider $\overline{\delta}\in(0,\delta)$ such that $\mathcal{L}^{d}(\Omega_{\overline{\delta}}\setminus\Omega_{\delta})<\rho$ and set $A_{\delta,\varepsilon}:=y^{-1}_{\varepsilon}(\Omega_{\delta})$. For $\varepsilon>0$ small enough, we have that $A_{\delta,\varepsilon}\subseteq\Omega_{\overline{\delta}}$. We rewrite
(3.15) $\int_{\Omega_{\delta}}|m_{\varepsilon}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x=\int_{\Omega_{\delta}\cap A_{\delta,\varepsilon}}|m_{\varepsilon}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x+\int_{\Omega_{\delta}\setminus A_{\delta,\varepsilon}}|m_{\varepsilon}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x.$
We notice that, by change of variable,
$\mathcal{L}^{d}(\Omega_{\delta})=\mathcal{L}^{d}(y_{\varepsilon}(A_{\delta,\varepsilon}))=\mathcal{L}^{d}(A_{\delta,\varepsilon})+\int_{A_{\delta,\varepsilon}}(\det\nabla y_{\varepsilon}-1){\,\rm d}x.$
Since $\det\nabla y_{\varepsilon}\to 1$ in $L^{1}(\Omega)$, $A_{\delta,\varepsilon}\subseteq\Omega_{\overline{\delta}}$, and $\mathcal{L}^{d}(\Omega_{\overline{\delta}}\setminus\Omega_{\delta})<\rho$, we conclude that
$\limsup_{\varepsilon\to 0}\mathcal{L}^{d}(\Omega_{\delta}\setminus A_{\delta,\varepsilon})$ $\leq\limsup_{\varepsilon\to 0}\mathcal{L}^{d}(\Omega_{\overline{\delta}})-\mathcal{L}^{d}(A_{\delta,\varepsilon})$
$=\limsup_{\varepsilon\to 0}\mathcal{L}^{d}(\Omega_{\overline{\delta}}\setminus\Omega_{\delta})-\int_{A_{\delta,\varepsilon}}(\det\nabla y_{\varepsilon}-1){\,\rm d}x<\rho.$
In turn, this implies that
(3.16) $\limsup_{\varepsilon\to 0}\int_{\Omega_{\delta}\setminus A_{\delta,\varepsilon}}|m_{\varepsilon}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x<\eta.$
To estimate the integral on $\Omega_{\delta}\cap A_{\delta,\varepsilon}$, we fix $\lambda\in(-1/p,0)$, extend $\mu_{\varepsilon}^{\delta}$ from $\Omega_{\delta}$ to a map $M_{\varepsilon}^{\delta}\in W^{1,2}(\mathbb{R}^{d};\mathbb{R}^{d})$ with $\|M_{\varepsilon}^{\delta}\|_{W^{1,2}(\mathbb{R}^{d})}\leq C(\Omega_{\delta})\|\mu_{\varepsilon}^{\delta}\|^{2}_{W^{1,2}(\Omega_{\delta})}$, and use the Lusin approximation of Sobolev functions: for every $\varepsilon>0$ there exists a set $G_{\delta,\varepsilon}$ such that $M_{\varepsilon}^{\delta}$ is $\varepsilon^{\lambda}$-Lipschitz on $G_{\delta,\varepsilon}$ and
(3.17) $\mathcal{L}^{d}(\mathbb{R}^{d}\setminus G_{\delta,\varepsilon})\leq\varepsilon^{-2\lambda}\int_{\mathbb{R}^{d}}|\nabla M_{\varepsilon}^{\delta}|^{2}{\,\rm d}x\leq C(\Omega_{\delta})\varepsilon^{-2\lambda}\|\mu_{\varepsilon}^{\delta}\|^{2}_{W^{1,2}(\Omega_{\delta})}.$
Setting $X_{\delta,\varepsilon}:=y_{\varepsilon}^{-1}(G_{\delta,\varepsilon})$ and noticing that $m_{\varepsilon}=M_{\varepsilon}^{\delta}$ on $\Omega_{\delta}\cap G_{\delta,\varepsilon}$, we have that
$\int_{\Omega_{\delta}\cap A_{\delta,\varepsilon}\cap X_{\delta,\varepsilon}\cap G_{\delta,\varepsilon}}|m_{\varepsilon}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x$
$\qquad=\int_{\Omega_{\delta}\cap A_{\delta,\varepsilon}\cap X_{\delta,\varepsilon}\cap G_{\delta,\varepsilon}}|M_{\varepsilon}^{\delta}\circ y_{\varepsilon}-M_{\varepsilon}^{\delta}|{\,\rm d}x\leq\varepsilon^{\lambda}\mathcal{L}^{d}(\Omega)\|y_{\varepsilon}-id\|_{C^{0}}\leq\varepsilon^{\frac{2}{p}+\lambda}\mathcal{L}^{d}(\Omega).$
Combining (3.17), the convergence rate (3.9), and a change of variable, we infer that
$\lim_{\varepsilon\to 0}\mathcal{L}^{d}(\Omega_{\delta}\setminus G_{\delta,\varepsilon})=0\qquad\text{and}\qquad\lim_{\varepsilon\to 0}\mathcal{L}^{d}(\Omega_{\delta}\setminus X_{\delta,\varepsilon})=0.$
Thus,
(3.18) $\limsup_{\varepsilon\to 0}\int_{\Omega_{\delta}\cap A_{\delta,\varepsilon}\setminus(X_{\delta,\varepsilon}\cap F_{\delta,\varepsilon})}|m_{\varepsilon}^{\delta}\circ y_{\varepsilon}-\mu_{\varepsilon}^{\delta}|{\,\rm d}x<\eta.$
Combining (3.14), (3.15), (3.16), and (3.18) yields
$\limsup_{\varepsilon\to 0}\int_{\Omega_{\delta}}|m_{\varepsilon}\circ y_{\varepsilon}-m|{\,\rm d}x<3\eta.$
Since $\eta>0$ is arbitrary, this implies that $m_{\varepsilon}\circ y_{\varepsilon}\to m$ in $L^{1}(\Omega_{\delta};\mathbb{R}^{d})$ for every $\delta>0$. By the equi-integrability of $m_{\varepsilon}\circ y_{\varepsilon}$, this entails the convergence in $L^{1}(\Omega;\mathbb{R}^{d})$. Furthermore, since strong $L^{1}$-convergence implies convergence a.e., we have $m_{\varepsilon}\circ y_{\varepsilon}\det\nabla y_{\varepsilon}\to m$ a.e. and $|m_{\varepsilon}\circ y_{\varepsilon}|\det\nabla y_{\varepsilon}=|m|=1$ a.e. in $\Omega$. Thus, $m\in W^{1,2}(\Omega;\mathbb{S}^{d-1})$ and $m_{\varepsilon}\circ y_{\varepsilon}\det\nabla y_{\varepsilon}\to m$ in $L^{r}(\Omega;\mathbb{R}^{d})$ for any $r\in[1,+\infty)$.
4. Proof of $\Gamma$-convergence
We are now able to prove Theorem 1.1.
Proof of Theorem 1.1.
We divide the proof in lower and upper bound for the $\Gamma$-convergence.
$\Gamma$-liminf inequality. We start with the lower bound. Let $(u_{\varepsilon},m_{\varepsilon})\in\mathcal{A}_{\varepsilon}^{w}$ be such that
$\sup_{\varepsilon>0}\,\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})<+\infty.$
By Proposition 3.1, there exist $u\in W^{1,2}_{w}(\Omega;\mathbb{R}^{d})$ and $m\in W^{1,2}(\Omega;\mathbb{S}^{d-1})$ such that, up to a subsequence, $u_{\varepsilon}\rightharpoonup u$ weakly in $W^{1,2}(\Omega;\mathbb{R}^{d})$, $m_{\varepsilon}\circ y_{\varepsilon}\,\det\nabla y_{\varepsilon}\to m$ in $L^{r}(\Omega;\mathbb{R}^{d})$ for any $1\leq r<\infty$, $m_{\varepsilon}\circ y_{\varepsilon}\to m$ in $L^{1}(\Omega;\mathbb{R}^{d})$, $\chi_{y_{\varepsilon}}\nabla m_{\varepsilon}\rightharpoonup\chi_{\Omega}\nabla m$ weakly in $L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$, and $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m$ in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$. Then, by Proposition 2.1 we have that
$\int_{\Omega}|\nabla m|^{2}{\,\rm d}x+\int_{\mathbb{R}^{d}}|\nabla v_{m}|^{2}{\,\rm d}x\leq\liminf_{\varepsilon\to 0}\int_{y_{\varepsilon}(\Omega)}|\nabla m_{\varepsilon}|^{2}{\,\rm d}x+\int_{\mathbb{R}^{d}}|\nabla v_{m_{\varepsilon}}|^{2}{\,\rm d}x.$
For a fixed $10<\alpha<$ 1 define the sets
$E_{\varepsilon}:=\{x\in\Omega:|\nabla u_{\varepsilon}(x)|<\varepsilon^{-\alpha}\}.$
Then, by the Chebyshev inequality
[…]
Thus, $\chi_{E_{\varepsilon}}\to 1$ in measure and $\chi_{E_{\varepsilon}}\nabla u_{\varepsilon}\rightharpoonup\nabla u$ in $L^{2}(\Omega;\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$. Indeed, for $21\leq r<$ 2 we may estimate
$\int_{\Omega\setminus E_{\varepsilon}}|\nabla u_{\varepsilon}-\nabla u|^{r}{\,\rm d}x$ $\leq\|\nabla u_{\varepsilon}-\nabla u\|^{\frac{1}{r}}_{L^{2}}\cdot\mathcal{L}^{d}(\Omega\setminus E_{\varepsilon})^{\frac{2-r}{2r}}\leq C\varepsilon^{\frac{(2-r)\alpha}{2r}}\to 0.$
Therefore, $\chi_{E_{\varepsilon}}\nabla u_{\varepsilon}=\nabla u_{\varepsilon}-\chi_{\Omega\setminus E_{\varepsilon}}\nabla u_{\varepsilon}\rightharpoonup\nabla u\text{ weakly in }L^{r}(\Omega;\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$. Since the weak limit in $L^{r}(\Omega;\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$ is unique, we get that $\chi_{E_{\varepsilon}}\nabla u\rightharpoonup\nabla u$ weakly in $L^{2}(\Omega;\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$
Moreover, since $m_{\varepsilon}\circ{y}_{\varepsilon}\det\nabla y_{\varepsilon}\to m$ strongly in $L^{r}(\Omega;\mathbb{R}^{d})$ for every $1\leq r<\infty$ then $(\det\nabla y_{\varepsilon})^{2}m_{\varepsilon}\circ y_{\varepsilon}\otimes m_{\varepsilon}\circ y_{\varepsilon}\to m\otimes m$ strongly in $L^{r}(\Omega;\mathbb{R}^{d})$. Thus, we have that $e(\nabla y_{\varepsilon},{m_{\varepsilon}\circ y_{\varepsilon}})\to e(m)$ and $\chi_{E_{\varepsilon}}e(\nabla y_{\varepsilon},{m_{\varepsilon}\circ y_{\varepsilon}})\to e(m)$ strongly in $L^{2}(\Omega;\mathbb{R}^{d})$.
Let us set
$\beta(\varepsilon):=(\exp(\varepsilon e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})-I-\varepsilon e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}))\nabla y_{\varepsilon}+\varepsilon^{2}e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})\nabla u_{\varepsilon}.$
Recalling that $\|e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})\|_{\infty}$ and $\|\exp(\varepsilon e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})\|_{\infty}$ are uniformly bounded and $\Phi(I)=D\Phi(I)=0$, by Taylor expansion we have that
(4.1) $\frac{1}{\varepsilon^{2}}$ $\int_{\Omega}W(I+\varepsilon\nabla u_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}){\,\rm d}x$
$=\frac{1}{\varepsilon^{2}}\int_{E_{\varepsilon}}W(I+\varepsilon\nabla u_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}){\,\rm d}x+\frac{1}{\varepsilon^{2}}\int_{\Omega\setminus E_{\varepsilon}}W(I+\varepsilon\nabla u_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}){\,\rm d}x$
$\geq\frac{1}{\varepsilon^{2}}\int_{E_{\varepsilon}}W(I+\varepsilon\nabla u_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}){\,\rm d}x=\frac{1}{\varepsilon^{2}}\int_{E_{\varepsilon}}\Phi(\exp(\varepsilon e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})(I+\varepsilon\nabla u_{\varepsilon})){\,\rm d}x$
$=\frac{1}{\varepsilon^{2}}\int_{E_{\varepsilon}}\Phi\big{(}I+\varepsilon(\nabla u_{\varepsilon}+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}))+\beta(\varepsilon)\big{)}{\,\rm d}x$
$\geq\frac{1}{{2}\varepsilon^{2}}\int_{E_{\varepsilon}}\mathbb{C}(\varepsilon(\bm{\epsilon}(u_{\varepsilon})+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}))+\beta(\varepsilon)):(\varepsilon(\bm{\epsilon}(u_{\varepsilon})+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}))+\beta(\varepsilon)){\,\rm d}x$
$\hskip 20.00003pt-\int_{E_{\varepsilon}}\frac{\eta(|\varepsilon(\nabla u_{\varepsilon}+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}))+\beta(\varepsilon)|)|\varepsilon(\nabla u_{\varepsilon}+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}))+\beta(\varepsilon)|^{2}}{\varepsilon^{2}}{\,\rm d}x$
$\geq\frac{1}{{2}}\int_{E_{\varepsilon}}\mathbb{C}(\bm{\epsilon}(u_{\varepsilon})+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})):(\bm{\epsilon}(u_{\varepsilon})+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})){\,\rm d}x$
$\hskip 20.00003pt-2\int_{E_{\varepsilon}}\eta(|\varepsilon(\nabla u_{\varepsilon}+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})+\beta(\varepsilon))|)(|\nabla u_{\varepsilon}+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})|^{2}+\varepsilon^{-2}|\beta(\varepsilon)|^{2}){\,\rm d}x$
$\hskip 20.00003pt-\frac{1}{2\varepsilon^{2}}\int_{E_{\varepsilon}}\mathbb{C}\beta(\varepsilon):\beta(\varepsilon)\,{\,\rm d}x+\frac{1}{\varepsilon}\int_{E_{\varepsilon}}\mathbb{C}(\bm{\epsilon}(u_{\varepsilon})+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})):\beta(\varepsilon){\,\rm d}x.$
Since $|\nabla u_{\varepsilon}|\leq\varepsilon^{-\alpha}$ in $E_{\varepsilon}$ and $\alpha\in(0,1)$, we have that $\|\varepsilon\nabla u_{\varepsilon}+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})\|_{L^{\infty}(E_{\varepsilon})}$ is bounded. Hence, we estimate
(4.2) $\|\beta(\varepsilon)\|_{L^{\infty}(E_{\varepsilon})}\leq C\varepsilon^{2}(1+\varepsilon^{-\alpha}).$
for some $C>0$ independent of $\varepsilon$. In particular, also $\|\beta(\varepsilon)\|_{L^{\infty}(E_{\varepsilon})}$ is bounded. Thus, up to a redefinition of $C$, we have that
(4.3) $\|\eta$ $(|\varepsilon(\nabla u_{\varepsilon}+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})+\beta(\varepsilon))|)\|_{L^{\infty}(E_{\varepsilon})}$
$\leq C\|\varepsilon(\nabla u_{\varepsilon}+e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})+\beta(\varepsilon))\|_{L^{\infty}(E_{\varepsilon})}\leq C\varepsilon^{1-\alpha}$
for $\varepsilon\in(0,1)$. Combining (4.2) and (4.3) and recalling that $e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon})\to e(m)$ in $L^{2}(\Omega;\mathbb{R}^{d})$ (see Proposition 3.1), passing to the liminf in (4.1) we infer that
$\liminf_{\varepsilon\to 0}\frac{1}{\varepsilon^{2}}\int_{\Omega}W(I$ $+\varepsilon\nabla u_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}){\,\rm d}x\geq\frac{1}{{2}}\int_{\Omega}\mathbb{C}(\bm{\epsilon}(u)+e(m)):(\bm{\epsilon}(u)+e(m)){\,\rm d}x.$
[…]
(4.4) $y_{\varepsilon}(x)=x+\varepsilon u_{0}(x),$
(4.5) $m_{\varepsilon}(x+\varepsilon u_{0}(x))\det\nabla y_{\varepsilon}(x)=m_{0}(x).$
We want to prove that
(4.6) $y_{\varepsilon}\to id\text{ in }W^{1,p}(\Omega;\mathbb{R}^{d}),$
(4.7) $u_{\varepsilon}=u_{0}\to u_{0}\text{ in }W^{1,2}(\Omega;\mathbb{R}^{d}),$
(4.8) $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m_{0}\text{ in }L^{2}(\mathbb{R}^{d};\mathbb{R}^{d}),$
(4.9) $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}\to\chi_{\Omega}\nabla m_{0}\text{ in }L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}),$
(4.10) $\limsup_{\varepsilon\to 0}\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})\leq\mathcal{G}_{0}(u_{0},m_{0}).$
First, we notice that $y_{\varepsilon}\to id$ in $W^{2,\infty}(\Omega;\mathbb{R}^{d})$, since $\|y_{\varepsilon}-id\|_{W^{2,\infty}}\leq\varepsilon\|u_{0}\|_{W^{2,\infty}}$. This implies the convergences (4.6) and (4.7).
The injectivity of $y_{\varepsilon}$ and the sign condition $\det\nabla y_{\varepsilon}>0$ follow from Theorem 2.2 ([8, Theorem 5.5-1]), since $\varepsilon\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\nabla u_{0}\|_{\infty}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq c(\Omega)$ for small enough $\varepsilon$.
Let us now notice that $y_{\varepsilon}$ is bi-Lipschitz for small enough $\varepsilon$. Indeed, there exist constants $L$ and $l>0$ (that may depend on $u_{0}$ but not on $x$ and $\varepsilon$), such that for small enough $\varepsilon$ it holds that $\sup_{x\in\Omega}|\nabla y_{\varepsilon}|=\sup_{x\in\Omega}|I+\varepsilon\nabla u_{0}(x)|=L_{\varepsilon}<L$ with $L_{\varepsilon}\to 1$ if $\varepsilon\to 0$. Moreover, $\inf_{x\in\Omega}\det\nabla y_{\varepsilon}=\inf_{x\in\Omega}\det(I+\varepsilon\nabla u_{0})=l_{\varepsilon}>l$ and $l_{\varepsilon}\to 1$ if $\varepsilon\to 0$. By the inverse function theorem $y_{\varepsilon}^{-1}$ is differentiable and $\nabla y_{\varepsilon}^{-1}(y_{\varepsilon}(x))=(\nabla y_{\varepsilon}(x))^{-1}\approx I-\varepsilon\nabla u_{0}(x)$. Thus, $m_{\varepsilon}$ defined in (4.5) can be explicitly written as
$m_{\varepsilon}(z)=\frac{1}{\det\nabla y_{\varepsilon}}m_{0}\circ y_{\varepsilon}^{-1}(z)\text{ for }z\in y_{\varepsilon}(\Omega).$
Hence, we have that
$|m_{\varepsilon}(z)|=\left|\frac{m_{0}(y_{\varepsilon}^{-1}(z))}{\det\nabla y_{\varepsilon}(y_{\varepsilon}^{-1}(z))}\right|\leq\frac{1}{l_{\varepsilon}}.$
By direct computation, we have that
$\nabla\left(\frac{m_{0}}{\det\nabla y_{\varepsilon}}\right)$ $=\frac{1}{\det\nabla y_{\varepsilon}}\nabla m_{0}+m_{0}\otimes\nabla\left(\frac{1}{\det\nabla y_{\varepsilon}}\right)$
$=\frac{1}{\det\nabla y_{\varepsilon}}\nabla m_{0}-\frac{1}{(\det\nabla y_{\varepsilon})^{2}}m_{0}\otimes\nabla(\det\nabla y_{\varepsilon}).$
This implies that
$\bigg{\|}\nabla\left(\frac{m_{0}}{\det\nabla y_{\varepsilon}}\right)\bigg{\|}_{L^{2}}$ $\leq\frac{\|\nabla m_{0}\|_{L^{2}}}{l_{\varepsilon}}+C\frac{\|\nabla\det\nabla y_{\varepsilon}\|_{L^{\infty}}}{l_{\varepsilon}},$
for some positive constant $C$ only depending on $\Omega$. This inequality, together with the fact that $\left|\nabla y^{-1}_{\varepsilon}(z)\right|\leq\frac{L_{\varepsilon}^{d-1}}{l_{\varepsilon}}$, provides
(4.11) $\int_{y_{\varepsilon}(\Omega)}|\nabla m_{\varepsilon}(z)|^{2}{\,\rm d}z$ $=\int_{y_{\varepsilon}(\Omega)}\left|\nabla\left(\frac{m_{0}(y^{-1}_{\varepsilon}(z))}{\det\nabla y_{\varepsilon}(y^{-1}_{\varepsilon}(z))}\right)\cdot\nabla y^{-1}_{\varepsilon}(z)\right|^{2}\det{\nabla y_{\varepsilon}}(y^{-1}_{\varepsilon}(z)){\,\rm d}z$
$\leq\int_{\Omega}\left|\nabla\left(\frac{m_{0}(x)}{\det\nabla y_{\varepsilon}(x)}\right)\right|^{2}\left|\nabla y^{-1}_{\varepsilon}(y_{\varepsilon}(x))\right|^{2}\det\nabla y_{\varepsilon}(x){\,\rm d}x$
$\leq\frac{2L_{\varepsilon}^{3d-2}}{l_{\varepsilon}^{4}}\|\nabla m_{0}(x)\|^{2}_{L^{2}}+C\frac{\|\nabla\det\nabla y_{\varepsilon}\|^{2}_{L^{\infty}}}{l_{\varepsilon}^{2}}.$
Thus, $m_{\varepsilon}\in W^{1,2}(y_{\varepsilon}(\Omega);\mathbb{R}^{d})$.
Let us define $\Omega_{\delta}:=\{x\in\Omega:\,\operatorname{dist}(x,\partial\Omega)>\delta\}$. In particular, $\Omega_{\delta}\subset\Omega\cap y_{\varepsilon}(\Omega)$ for $\varepsilon>0$ small enough, due to the uniform convergence of $y_{\varepsilon}$ to the identity. For such $\varepsilon$ and for $z\in\Omega_{\delta}$ we estimate
$|m_{\varepsilon}(z)-m_{0}(z)|$ $\leq\left|\frac{m_{0}(y_{\varepsilon}^{-1}(z))}{\det\nabla y_{\varepsilon}(y_{\varepsilon}^{-1}(z))}-\frac{m_{0}(z)}{\det\nabla y_{\varepsilon}(y_{\varepsilon}^{-1}(z))}\right|+\left|\frac{m_{0}(z)}{\det\nabla y_{\varepsilon}(y_{\varepsilon}^{-1}(z))}-m_{0}(z)\right|$
$\leq\frac{1}{\det\nabla y_{\varepsilon}(y_{\varepsilon}^{-1}(z))}\left|m_{0}(y_{\varepsilon}^{-1}(z))-m_{0}(z)\right|+|m_{0}(z)|\cdot\left|\frac{1}{\det\nabla y_{\varepsilon}(y_{\varepsilon}^{-1}(z))}-1\right|.$
Therefore, taking the squares, integrating over $\Omega_{\delta}$, and letting $\varepsilon\to 0$ we infer that $m_{\varepsilon}\to m_{0}$ in $L^{2}(\Omega_{\delta};\mathbb{R}^{d})$, for every $\delta>0$. Arguing as in the proof of Proposition 3.1, it is not difficult to see that $\chi_{y_{\varepsilon}(\Omega)}m_{\varepsilon}\to\chi_{\Omega}m_{0}$ strongly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$ and $\chi_{y_{\varepsilon}(\Omega)}\nabla m_{\varepsilon}\to\chi_{\Omega}\nabla m_{0}$ weakly in $L^{2}(\mathbb{R}^{d};\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}^{d\times d}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})$. Moreover, by (4.11) we can estimate the exchange energy as
$\limsup_{\varepsilon\to 0}\int_{y_{\varepsilon}(\Omega)}|\nabla m_{\varepsilon}(y)|^{2}{\,\rm d}y$ $\leq\limsup_{\varepsilon\to 0}\int_{\Omega}\frac{L_{\varepsilon}^{3d-2}}{l_{\varepsilon}^{4}}|\nabla m_{0}(x)|^{2}{\,\rm d}x+C\frac{\|\nabla\det\nabla y_{\varepsilon}\|^{2}_{L^{\infty}}}{l_{\varepsilon}^{2}}=\int_{\Omega}|\nabla m_{0}(x)|^{2}{\,\rm d}x.$
For a.e. $x\in\Omega$, the elastic energy has the following form by Taylor’s expansion and by definition of $m_{\varepsilon}$
$\frac{1}{\varepsilon^{2}}W(I+\varepsilon\nabla u_{0},m_{\varepsilon}\circ y_{\varepsilon})$ $=\frac{1}{\varepsilon^{2}}\Phi(\exp(\varepsilon e(\nabla y_{\varepsilon},m_{\varepsilon}\circ y_{\varepsilon}))(I+\varepsilon\nabla u_{0}))$
$=\frac{1}{\varepsilon^{2}}\Phi(\exp(\varepsilon e(m_{0}))(I+\varepsilon\nabla u_{0}))$
$=\frac{1}{\varepsilon^{2}}\Phi(I+\varepsilon(\nabla u_{0}+e(m_{0}))+o(\varepsilon))$
$=\frac{1}{{2}}\mathbb{C}(\bm{\epsilon}(u_{0})+e(m_{0})):(\bm{\epsilon}(u_{0})+e(m_{0}))+O(1)$
with $O(1)\to 0$ as $\varepsilon\to 0$ uniformly in $\Omega$. Then (d) and the Dominated Convergence Theorem give
$\limsup\limits_{\varepsilon\to 0}\frac{1}{\varepsilon^{2}}\int_{\Omega}W(I+\varepsilon\nabla u_{0},m_{\varepsilon}\circ y_{\varepsilon}){\,\rm d}x$ $=\frac{1}{{2}}\int_{\Omega}\mathbb{C}(\bm{\epsilon}(u_{0})+e(m_{0})):(\bm{\epsilon}(u_{0})+e(m_{0})){\,\rm d}x.$
We conclude that
$\limsup_{\varepsilon\to 0}\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})$ $=\limsup_{\varepsilon\to 0}\frac{1}{\varepsilon^{2}}\int_{\Omega}W(I+\varepsilon\nabla u_{0},m_{0}){\,\rm d}x+\frac{1}{2}\int_{y_{\varepsilon}(\Omega)}|\nabla m_{\varepsilon}|^{2}{\,\rm d}y+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m_{\varepsilon}}|^{2}{\,\rm d}z$
$\leq\frac{1}{{2}}\int_{\Omega}\mathbb{C}(\bm{\epsilon}(u_{0})+e(m_{0})):(\bm{\epsilon}(u_{0})+e(m_{0})){\,\rm d}x+\frac{1}{2}\int_{\Omega}|\nabla m_{0}|^{2}{\,\rm d}x+\frac{\mu_{0}}{2}\int_{\mathbb{R}^{d}}|\nabla v_{m_{0}}|^{2}{\,\rm d}z$
$=\mathcal{G}_{0}(u_{0},m_{0}),$
where, for the convergence of $\nabla v_{m_{\varepsilon}}$, we have used Proposition 2.1.
For a general $u_{0}\in W^{1,2}_{w}(\Omega;\mathbb{R}^{d})$ we conclude by a standard diagonal argument. ∎
We conclude this section with the proof of convergence of minimizers of $\mathcal{F}_{\varepsilon}$ to minimizers of $\mathcal{F}_{0}$, namely, Theorems 1.2 and 1.3. We start with the coercivity of $\mathcal{F}_{\varepsilon}$.
Proposition 4.1.
Let $(u_{\varepsilon},m_{\varepsilon})\in\mathcal{A}^{w}_{\varepsilon}$ and $K>0$ be such that
$\sup_{\varepsilon>0}\mathcal{F}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})\leq K.$
Then, there exists $C(K)>0$ such that
$\sup_{\varepsilon>0}\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})\leq C(K).$
Proof.
Recall that $\mathcal{F}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon}):=\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})-\mathscr{L}(u_{\varepsilon})-\mathscr{M}(y_{\varepsilon},m_{\varepsilon})$, with $|\mathscr{L}(u_{\varepsilon})|\leq C_{L}\|u_{\varepsilon}\|_{L^{2}(\Omega;\mathbb{R}^{d})}$ and $|\mathscr{M}(y_{\varepsilon},m_{\varepsilon})|\leq C_{M}\|m_{\varepsilon}\|_{L^{2}(y_{\varepsilon}(\Omega);\mathbb{R}^{d})}$. Moreover, by Lemma 3.3 we have
$\|u_{\varepsilon}\|_{L^{2}(\Omega;\mathbb{R}^{d})}^{2}\leq C\bigg{(}\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})+\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\int_{\Gamma}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}\,{\,\rm d}\mathcal{H}^{d-1}+1\bigg{)},$
while arguing as in (3.10) we deduce from the assumption (e) that
$\|m_{\varepsilon}\|_{L^{2}(y_{\varepsilon}(\Omega);\mathbb{R}^{d})}^{2}=\int_{\Omega}\frac{1}{\det\nabla y_{\varepsilon}}{\,\rm d}x\leq C(\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})+1).$
Summing up the estimates above, we obtain
$\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})=\mathcal{F}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})+\mathscr{L}(u_{\varepsilon})+\mathscr{M}(y_{\varepsilon},m_{\varepsilon})\leq K+C\sqrt{\mathcal{G}_{\varepsilon}(u_{\varepsilon},m_{\varepsilon})}+C,$
This concludes the proof of the proposition. ∎
We are now able to prove Theorem 1.2.
Proof of Theorem 1.2.
Consider a sequence $\varepsilon_{k}\to 0$ and recall that
$s_{\varepsilon}:=\inf\{\mathcal{F}_{\varepsilon}(u,m):(u,m)\in\mathcal{A}^{w}_{\varepsilon}\}\qquad\text{ and }\qquad s_{0}:=\inf\{\mathcal{F}_{0}(u,m):(u,m)\in\mathcal{A}^{w}_{0}\}.$
It is standard to show that $\mathcal{F}_{0}$ has a minimizer $(u_{0},m_{0})\in W^{1,2}_{w}(\Omega;\mathbb{R}^{d})\times W^{1,2}(\Omega;\mathbb{S}^{d-1})$ on $\mathcal{A}_{0}^{w}$. It is also straightforward to check that $\inf\mathcal{F}_{\varepsilon_{k}}(u,m)$ is bounded w.r.t. $\varepsilon$. Since the functionals $\mathscr{L}$ and $\mathscr{M}$ do not depend on $\varepsilon$, from $\Gamma$-convergence of $\mathcal{G}_{\varepsilon}$ we deduce that $s_{\varepsilon_{k}}\to s_{0}$ and thus $\mathcal{F}_{\varepsilon_{k}}(u_{\varepsilon_{k}},m_{\varepsilon_{k}})\to\mathcal{F}_{0}(u_{0},m_{0})$. Propositions 4.1 and 3.1 imply that $u_{\varepsilon_{k}}\rightharpoonup u_{0}$ weakly in $W^{1,2}_{w}(\Omega;\mathbb{R}^{d})$, $\chi_{y_{\varepsilon_{k}(\Omega)}}m_{\varepsilon_{k}}\to\chi_{\Omega}m_{0}$ strongly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and $\chi_{y_{\varepsilon_{k}(\Omega)}}\nabla m_{\varepsilon_{k}}\rightharpoonup\chi_{\Omega}\nabla m_{0}$ weakly in $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, and $\mathcal{G}_{\varepsilon_{k}}(u_{\varepsilon_{k}},m_{\varepsilon_{k}})\to\mathcal{G}(u_{0},m_{0})$. ∎
Proof of Theorem 1.3.
The convergence of the sequence $u_{\varepsilon}$ can be improved to a strong convergence in $W^{1,2}(\Omega;\mathbb{R}^{d})$ arguing as in [16, Subsection 7.2] (see also [6, Subsection 3.2]).
Acknowledgements
The work of S.A. was funded by the Austrian Science Fund (FWF) through the projects P 35359-N and ESP 65, by the Italian Ministry of Education and Research through the PRIN 2022 project “Variational Analysis of Complex Systems in Material Science, Physics and Biology” No. 2022HKBF5C, and by the University of Naples Federico II through the FRA project “Regularity and Singularity in Analysis, PDEs, and Applied Sciences”. S.A. also acknowledges the hospitality of the University of Vienna and the TU Wien, where part of this work has been done. M.K. was supported by the GAČR-FWF project 21-06569K and by the Ministry of Education, Youth and Sport of the Czech Republic through the bilateral project 8J22AT017. A.M. was supported by the European Unions Horizon 2020 research and innovation programme under the Marie Składowska-Curie grant agreement No. 847693.
Finally, S.A. and A.M. acknowledge the kind hospitality of the ESI Institute (Vienna) during the workshop “New perspective on Shape and Topology Optimization”.